# DeepTrace: Learning to Optimize Contact Tracing in Epidemic Networks with Graph Neural Networks

### v3

@@ Abstract
Digital contact tracing aims to curb epidemics by identifying and mitigating public health emergencies through technology. Backward contact tracing, which tracks the sources of infection, proved crucial in places like Japan for identifying COVID-19 infections from superspreading events. This paper presents a novel perspective of digital contact tracing as online graph exploration and addresses the forward and backward contact tracing problem as a maximum-likelihood (ML) estimation problem using iterative epidemic network data sampling. The challenge lies in the combinatorial complexity and rapid spread of infections. We introduce DeepTrace, an algorithm based on a Graph Neural Network (GNN) that iteratively updates its estimations as new contact tracing data is collected, learning to optimize the maximum likelihood estimation by utilizing topological features to accelerate learning and improve convergence. The contact tracing process combines either BFS or DFS to expand the network and trace the infection source, ensuring comprehensive and efficient exploration. Additionally, the GNN model is fine-tuned through a two-phase approach: pre-training with synthetic networks to approximate likelihood probabilities and fine-tuning with high-quality data to refine the model. Using COVID-19 variant data, we illustrate that DeepTrace surpasses current methods in identifying superspreaders, providing a robust basis for a scalable digital contact tracing strategy.
@@ I Introduction
Since the onset of the COVID-19 pandemic, a multitude of mobile applications [1, 2, 3, 4] have been deployed to detect individuals potentially exposed to the SARS-CoV-2 coronavirus and mitigate its transmission. These applications offer exposure notifications as an adjunct to conventional contact tracing methodologies, which are often resource-intensive and time-consuming. Utilizing network science-based data analytics and machine learning, these solutions show promise in automating widespread coverage, crucial for monitoring the rapid spread of infectious diseases and their variants [5]. The goal of digital contact tracing is thus to diminish the spread of the epidemic, but this is complicated by factors such as the speed of disease spreading, asymptomatic infections, and the infectiousness characteristics, i.e., a few superspreaders affecting the whole (e.g., approximately $20\%$ of infected individuals responsible for $80\%$ of transmissions [6]). Designing robust digital contact tracing automated by network science and machine learning is still in its infancy.
It has been recently shown that [7, 8, 9, 2] compared with ‘forward’ tracing (i.e., to whom disease spreads), ‘backward’ tracing (from whom disease spreads) can be more effective due to overlooked biases arising from the heterogeneity in contacts. Tracing sources of spreading (i.e., backward contact tracing), as had been used in Japan [8], has proven effective as going backward can pick up infections that might otherwise be missed at superspreading events. The ability to identify the source of spreading (i.e., the Patient Zero or superspreaders in a pandemic) is essential to contact tracing. In this paper, we formulate a jointly forward and backward contact tracing problem that leverages the maximum likelihood estimation as a subproblem to design contact tracing algorithms based on a network centrality design approach.
@@ I-A Related Works
The contagion source inference problem was first studied in the seminal work [10] as a maximum likelihood estimation problem based on a Susceptible-Infectious (SI) model [11], which is a special case of the classical susceptible-infected-recover (SIR) model in epidemiology [12, 13, 14]. This problem becomes challenging when dealing with a very large graph, as the sheer scale amplifies the intricacy of tracking numerous spreading likelihoods across its vast network, compounding the complexity of calculating permissible permutations for each node. For the degree-regular tree special case, this problem can be solved in polynomial time by a network centrality approach, i.e., the rumor centrality [10, 15], which is proportional to the number of permitted permutations. Equivalently, this optimal likelihood estimate is the graph centroid [16, 17, 18]. The network centrality approach [10] has been extended to scenarios like random trees [19, 20], multiple snapshot observations [21], multiple sources [22] and the epidemic centrality for a disease pandemic in [23]. Unlike the snapshot model [10], other studies [24, 25, 26, 27] use time series data, Bayesian analysis and stochastic processes to solve the contagion source inference problem via sequential and quickest detection techniques. For a comprehensive overview of the work on the detection of contagion sources, we refer the reader to [18].
@@
However, computing the ML estimator for the source on large epidemic networks can be computationally challenging [10, 23], particularly as the network size grows and complexity increases with the presence of cycles. Thus, using Graph Neural Networks (GNNs) becomes a viable option. Various studies have used GNNs for pandemic control [28, 29, 30, 31]. Gatta [28] focuses on estimating epidemiological model parameters, differing from other works [29, 30, 31] that aim to identify patient zero (superspreaders). Shah [29] examines SIR and SEIR models with graph convolutional networks, but these are limited to “transductive" settings, unable to handle graph inputs of varying sizes. On the other hand, the work in [30] also considers the same discrete time SIR spreading model, but in the training dataset of this paper, susceptible nodes were pre-assigned values such that susceptible nodes are unlikely to be the source estimator using external expert knowledge. For both works [30, 29], the label of each piece of data in the training set is the source of the outbreak but not the ML estimator of the source according to their spreading model. Thus, the labels used in previous works do not have statistical significance; sometimes, the source may not equal the ML estimator, which introduces noise into the training data.
@@
To address the issues related to the training set [30, 29], and the computational complexity challenges discussed in [15, 23], we propose a GNN-based framework with a two-phase semi-supervised training approach. In this framework, we use graphs that contain the ML estimator for the superspreader during the fine-tuning phase. Specifically, we adopt the SI spreading model, as thoroughly studied in [15, 23, 18]. It has been established that the ML estimator is equivalent to certain network centers [17, 10, 23]. This equivalence allows us to easily obtain high-quality labeled data for specific network topologies, thereby improving the accuracy of identifying the most likely superspreader.
@@ I-B Main Contributions
  • We frame digital contact tracing as an online graph exploration problem, where backward contact tracing is solved as a maximum likelihood estimation problem that captures the dynamic effects of the pandemic’s spread over time. This iterative process involves a challenging nonconvex optimization problem. We tackle this using modified graph neural networks to predict network centrality measures. Applied to epidemic networks of varying sizes, our low-complexity algorithm, DeepTrace, progressively refines the prediction of the most likely source of contagion as more data becomes available.
@@
  • Using data sets of both synthetic epidemic networks and contact tracing data during the COVID-19 pandemic in Taiwan and Hong Kong, we demonstrate that Algorithm DeepTrace outperforms the state-of-the-art heuristics in [10, 23].
@@ II Contact Tracing Problem Formulation
Let us first introduce the spreading models including one relevant to COVID-$19$ in [14] and then present the digital contact tracing optimization problem formulation followed by its mathematical analysis.
@@ II-A Epidemic Spreading Model
There are numerous models in the literature in the study of the spread of infectious diseases ranging from commonly well-known ones like the Susceptible-Infectious (SI) model to the Susceptible-Infectious-Recovery-Susceptible (SIRS) models [12, 24, 13, 14]. Despite being the most basic one, the SI model can accurately model a pandemic in its early stage (e.g., COVID-$19$ ca. 2020), whereas other SIRS models [14] incorporate complex factors like virus mutation, societal response (e.g., mask wearing and social distancing) and the use of vaccines (e.g., COVID-$19$ ca. 2021).
We focus primarily on the SI model for analysis and use the more accurate COVID-$19$ model [14] for numerical performance evaluation. The SI model assumes that there are only two types of individuals, say susceptible and infectious during the epidemic spreading. Also, we follow the assumption that the times between infection events are independent and exponentially distributed [15] with an average infection time (with most estimates of COVID-$19$ incubation period ranging from 1 to 14 days [14]). The virus spreads in a social contact network cab be modeled by a diffusion process on a graph denoted by $\mathbb{G}$. Let $V(\mathbb{G})$ and $E(\mathbb{G})$ denote the set of nodes and edges in $\mathbb{G}$, which represent people and their social contacts (an edge exists between two persons if, say, they are within 6 feet for more than 15 minutes) respectively. We shall call $\mathbb{G}$ the underlying network, where the topology of $\mathbb{G}$ poses a constraint to the epidemic spreading [32, 15, 23]. The spreading of the virus starts from a node in $\mathbb{G}$ (i.e., the superspreader) and spreads to other nodes under the constraint that a susceptible node can be infected only if at least one of its neighboring nodes has been infected. At some point in time, all infected nodes in $\mathbb{G}$ form a connected subgraph of $\mathbb{G}$, denoted as $\mathbb{G}_{N}$ and is called the epidemic network, where $N$ represents the number of nodes in the epidemic network.
@@ II-B Forward and Backward Contact Tracing
At various stages of contact tracing, we model an instantaneous snapshot of a subgraph of $\mathbb{G}_{N}$ that we call the contact tracing network. This contact tracing network grows by one (infected) node at each stage. If the traced node is infected, we continue to trace the node’s neighbors; otherwise, we stop tracing along a path involving this node. Let $G_{n}$ denote the contact tracing network with $n$ nodes at the $n$-th stage of tracing (thus the index case is $G_{1}$), hence we have $G_{n}\subseteq\mathbb{G}_{N}\subseteq\mathbb{G}$. A goal of contact tracing is to find the node in $\mathbb{G}_{N}$ that is most likely to be the superspreader, as shown in Fig. 1. However, this most-likely superspreader may not yet be in $G_{n}$, meaning that the contact tracing effort is still in its early stage or not fast enough (relative to the pandemic spreading speed). In such a case, backward contact tracing should yield an estimate as close as possible to this most-likely superspreader. In other words, given the available data at the $n$-th stage, the contact tracer finds the node in $G_{n}$ that is the fewest number of hops away from the most-likely superspreader in $\mathbb{G}_{N}$ (i.e., the optimal maximum likelihood estimate had this $\mathbb{G}_{N}$ been given entirely upfront to the contact tracer [10]).
@@
Given the data $G_{n}$ harvested by contact tracing at the $n$-th stage, we have the following maximum-likelihood estimation (MLE) problem:
@@
where $\mathbb{P}(G_{n}\mid v)$ is the likelihood function and $\hat{v}$ is the most-likely superspreader of the current observed $G_{n}$. The key challenge is that $\mathbb{G}_{N}$ is unknown to the contact tracer who has to consider:
@@
Answering both the forward and backward contact tracing jointly constitutes an iterative statistical inference process to track the most likely superspreader in the entire epidemic network. Specifically, for the forward contact tracing subproblem, it is natural to enlarge the forward contact tracing network with breadth-first search (BFS) or depth-first search (DFS) graph traversal starting from the index case. For the backward contact tracing subproblem, let us suppose that a given node $v$ in $G_{n}$ is the superspreader. Then, starting from this particular node, there are a number of possible ways to infect all the other nodes consistent with the given graph $G_{n}$ harvested by contact tracing at the $n$-th stage. Even though $\mathbb{G}_{N}$ is unknown, it is reasonable for a contact tracer to assume a possible infection order in $G_{n}$ by a permitted permutation given as $\sigma=\{v_{1}=v,v_{2},...,v_{n}\}$. Thus, we can reformulate problem (1) as:
@@
where $\Omega(G_{n}\mid v)$ is the collection of all permitted permutations for $G_{n}$ rooted at $v$, and $\mathbb{P}(\sigma|v)$ is the probability of the permitted permutation $\sigma$ with $v$ as source. Under the SI spreading model used in [15, 10, 23], the probability of a susceptible node being infected in the next time period is uniformly distributed among all outreaching edges from the infected nodes at the boundary of $G_{n}$, and the probability of a permitted permutation is as follows:
@@ II-C Regular Trees and General Graphs
From (3), we can observe that when $\mathbb{G}$ is a regular tree, both $\Phi_{i}$ and $d(v_{i})$ are constants which implies the value of (3) is also a constant. Thus, the computation of the likelihood only depends on the value of $|\Omega(G_{n},v)|$ in (2), which simplifies the computation. On the other hand, it is still unknown whether there exists an efficient method for calculating the likelihood when $\mathbb{G}$ is a general graph since $\mathbb{P}(\sigma\mid v)$ is different for each possible $\sigma\in\Omega(G_{n},v)$ and the size of $\Omega(G_{n},v)$ grows exponentially with $n$. To overcome this computational complexity issue, our goal is to leverage GNN to find the MLE. Various theoretical results on regular trees or unicyclic regular graphs [10, 23] provide us a way to generate high quality dataset, the label of each piece of data is the exact MLE of the epidemic graph, for training GNN. Moreover, the analysis in Theorem 1, 2 shows the best and the worst possible outcomes during the contact tracing on regular trees, which provides us reasonable bounds on the performance of the contact tracing on general graphs.
Coming back to the contact tracer starts from the index case $G_{1}$ and collects more data in a forward manner (i.e., enlarging $G_{n}$ in (2)), the contact tracer also predicts the superspreader for that instant by solving (1). Intuitively, this means that as the contact tracing subgraph $G_{n}$ grows, the contact tracer desires this prediction to be closer (in terms of the number of hops in $\mathbb{G}_{N}$) to the most likely superspreader in $\mathbb{G}_{N}$. Taking this into account, how should $G_{n}$ grow in size in forward contact tracing? We propose to enlarge $G_{n}$ using the BFS and the DFS graph traversal algorithms [33, 34], which is described in detail in the next section.

### v4

@@ Abstract
Digital contact tracing aims to curb epidemics by identifying and mitigating public health emergencies through technology. Backward contact tracing, which tracks the sources of infection, proved crucial in places like Japan for identifying COVID-19 infections from superspreading events. This paper presents a novel perspective on digital contact tracing by modeling it as an online graph exploration problem, framing forward and backward tracing strategies as maximum-likelihood estimation tasks that leverage iterative sampling of epidemic network data. The challenge lies in the combinatorial complexity and rapid spread of infections. We introduce DeepTrace, an algorithm based on a Graph Neural Network that iteratively updates its estimations as new contact tracing data is collected, learning to optimize the maximum likelihood estimation by utilizing topological features to accelerate learning and improve convergence. The contact tracing process combines either BFS or DFS to expand the network and trace the infection source, ensuring efficient real-time exploration. Additionally, the GNN model is fine-tuned through a two-phase approach: pre-training with synthetic networks to approximate likelihood probabilities and fine-tuning with high-quality data to refine the model. Using COVID-19 variant data, we illustrate that DeepTrace surpasses current methods in identifying superspreaders, providing a robust basis for a scalable digital contact tracing strategy.
@@ I Introduction
During the early days of the COVID-19 pandemic, a multitude of mobile applications [1, 2, 3, 4] were deployed to detect individuals potentially exposed to the SARS-CoV-2 coronavirus and mitigate its transmission. Digital contact tracing, driven by network analytics and machine learning, continues to play a critical role in epidemic surveillance and preventing hospital-acquired infections by optimizing coverage and tracking the spread of infectious diseases [5]. Its primary objective is to curb the transmission of epidemics, but this is complicated by challenges such as rapid disease spread, asymptomatic carriers, and the disproportionate impact of superspreaders (e.g., approximately $20\%$ of infected individuals account for $80\%$ of transmissions [6]). The development of robust, automated digital contact tracing systems leveraging network science and machine learning remains in its early stages.
It has been recently shown in [7, 8, 9, 2] that compared with ‘forward’ tracing (i.e., to whom disease spreads), ‘backward’ tracing (from whom disease spreads) can be more effective due to overlooked biases arising from the heterogeneity in contacts. Tracing sources of spreading (i.e., backward contact tracing), as had been used in Japan [8], has proven effective as going backward can pick up infections that might otherwise be missed at superspreading events. The ability to identify the source of spreading (i.e., the Patient Zero or superspreaders in a pandemic) is essential to contact tracing. In this paper, we propose a novel formulation of the contact tracing problem that integrates both forward and backward tracing, utilizing maximum likelihood estimation as a subproblem. Our approach focuses on optimizing who to track and how to track through graph neural network learning, employing network centrality metrics to design fast and efficient contact tracing algorithms.
@@ I-A Related Works
The contagion source inference problem was first studied in the seminal work [10] as a maximum likelihood estimation problem based on a Susceptible-Infectious (SI) model [11], which is a special case of the classical Susceptible-Infected-Recover (SIR) model in epidemiology [12, 13, 14]. This problem becomes challenging with large graphs, as the scale complicates tracking spreading likelihoods and calculating permissible permutations for each node. However, the degree-regular tree special case can be solved in polynomial time using a network centrality approach known as rumor centrality [10, 15], which is proportional to the number of permitted permutations. Equivalently, this optimal likelihood estimate is the tree centroid [16, 17, 18]. The network centrality approach [10] has been extended to scenarios like random trees [19, 20], multiple snapshot observations [21], multiple sources [22] and the epidemic centrality for a disease pandemic in [23]. Unlike the snapshot model [10], other studies [24, 25, 26, 27] use time series data, Bayesian analysis, and stochastic processes to solve the contagion source inference problem via sequential and quickest detection techniques. For a comprehensive overview of the work on detecting contagion sources, we refer the reader to [18].
@@
However, computing the ML estimator for the source on large epidemic networks can be computationally challenging [10, 23], particularly as the network size grows and complexity increases with the presence of cycles. Thus, using Graph Neural Networks (GNNs) is a viable option. Various studies have used GNNs for pandemic control [28, 29, 30, 31]. Gatta [28] focuses on estimating epidemiological model parameters, differing from other works [29, 30, 31] that aim to identify patient zero (superspreaders). Shah [29] examines SIR and SEIR models with graph convolutional networks, but these are limited to transductive settings and unable to handle graph inputs of varying sizes. In [30], susceptible nodes were pre-labeled as unlikely sources. Meanwhile, [30, 29] used the true outbreak source as the label, which can diverge from the maximum-likelihood (ML) estimator. Using the ML estimator as the label, we align training with identifying the node most likely to explain the data, thus reducing bias and improving accuracy.
@@
To address the issues related to the training set [30, 29], and the computational complexity challenges discussed in [15, 23], we propose a GNN-based framework with a two-phase semi-supervised training approach. In this framework, we use graph data with the ML estimator as the superspreader during the fine-tuning phase. Specifically, we adopt the SI spreading model, as thoroughly studied in [15, 23, 18]. It has been established that the ML estimator is equivalent to certain network centers [17, 10, 23]. This equivalence allows us to easily obtain high-quality labeled data for specific network topologies, thereby improving the accuracy of identifying the most likely superspreader.
@@ I-B Main Contributions
  • We frame digital contact tracing as an online graph exploration problem, deriving both forward and backward tracing strategies from a maximum likelihood estimation that captures the pandemic’s dynamic spread. To address the challenging nonconvex optimization involved, we employ modified graph neural networks to predict network centrality measures. Our low-complexity algorithm, DeepTrace, refines the prediction of the most likely contagion source as data is collected, effectively applying it to epidemic networks of various sizes.
@@
  • Using datasets from synthetic epidemic networks and COVID-19 contact tracing in Taiwan and Hong Kong, we show that Algorithm DeepTrace outperforms state-of-the-art heuristics from [10, 23].
@@ II Contact Tracing Problem Formulation
Let us first introduce the spreading models, including one relevant to COVID-$19$ in [14], and then present the digital contact tracing optimization problem formulation followed by its mathematical analysis.
@@ II-A Epidemic Spreading Model
There are numerous models in the literature in the study of the spread of infectious diseases ranging from commonly well-known ones like the Susceptible-Infectious model to the Susceptible-Infectious-Recovery-Susceptible (SIRS) models [12, 24, 13, 14]. Despite being the most basic one, the SI model can accurately model a pandemic in its early stage (e.g., COVID-$19$ ca. 2020), whereas other SIRS models [14] incorporate complex factors like virus mutation, societal response (e.g., mask-wearing and social distancing) and the use of vaccines (e.g., COVID-$19$ ca. 2021).
We primarily focus on the SI model for our analysis, as the characteristics of the COVID-19 virus can persist in individuals for extended periods without a clear recovery or removal phase. Studies have shown that SARS-CoV-2 can remain detectable in patients for several weeks, even after symptoms subside [32]. This persistent detectability supports using the SI model in the early stage of the outbreak. The SI model provides a simplified yet relevant framework for our analytical purposes, focusing on the essential aspects of identifying the initial source of infection. The SI model assumes that there are only two types of individuals, susceptible and infectious, during the epidemic spreading. Also, we follow the assumption that the times between infection events are independent and exponentially distributed [15] with an average infection time (with most estimates of COVID-$19$ incubation period ranging from 1 to 14 days [14]). The virus spreads in a social contact network can be modeled by a diffusion process on a graph denoted by $\mathbb{G}$. Let $V(\mathbb{G})$ and $E(\mathbb{G})$ denote the set of nodes and edges in $\mathbb{G}$, which represent people and their social contacts (an edge exists between two persons if, say, they are within 6 feet for more than 15 minutes) respectively. We shall call $\mathbb{G}$ the underlying network, where the topology of $\mathbb{G}$ poses a constraint to the epidemic spreading [33, 15, 23]. The spreading of the virus starts from a node in $\mathbb{G}$ (i.e., the superspreader) and spreads to other nodes under the constraint that a susceptible node can be infected only if at least one of its neighboring nodes has been infected. At some point in time, all infected nodes in $\mathbb{G}$ form a connected subgraph of $\mathbb{G}$, denoted as $\mathbb{G}_{N}$ and is called the epidemic network, where $N$ represents the number of nodes in the epidemic network.
@@ II-B Forward and Backward Contact Tracing
At various stages of contact tracing, we model an instantaneous snapshot of a subgraph of $\mathbb{G}_{N}$ that we call the contact tracing network. This contact tracing network grows by one (infected) node at each stage. If the traced node is infected, we continue to trace the node’s neighbors; otherwise, we stop tracing along a path involving this node. Let $G_{n}$ denote the contact tracing network with $n$ nodes at the $n$-th stage of tracing (thus, the index case is $G_{1}$). Hence we have $G_{n}\subseteq\mathbb{G}_{N}\subseteq\mathbb{G}$. A goal of contact tracing is to find the node in $\mathbb{G}_{N}$ that is most likely to be the superspreader, as shown in Fig. 1. However, this most likely superspreader may not yet be in $G_{n}$, meaning that the contact tracing effort is still in its early stage or not fast enough (relative to the pandemic spreading speed). In such a case, backward contact tracing should yield an estimate as close as possible to this most likely superspreader. In other words, given the available data at the $n$-th stage, the contact tracer finds the node in $G_{n}$ that is the fewest number of hops away from the most likely superspreader in $\mathbb{G}_{N}$ (i.e., the optimal maximum likelihood estimate had this $\mathbb{G}_{N}$ been given entirely upfront to the contact tracer [10]).
@@
Given the data $G_{n}$ harvested by contact tracing at the $n$-th stage, we have the following maximum-likelihood estimation problem:
@@
where $\mathbb{P}(G_{n}\mid v)$ is the likelihood function and $\hat{v}$ is the most-likely superspreader of the current observed $G_{n}$. The key challenge is that $\mathbb{G}_{N}$ is unknown to the contact tracer, who has to consider:
@@
Answering both the forward and backward contact tracing jointly constitutes an iterative statistical inference process to track the most likely superspreader in the epidemic network. Specifically, for the forward contact tracing subproblem, it is natural to enlarge the forward contact tracing network with breadth-first search (BFS) or depth-first search (DFS) graph traversal starting from the index case. For the backward contact tracing subproblem, suppose that a given node $v$ in $G_{n}$ is the superspreader. Then, starting from this particular node, there are many possible ways to infect all the other nodes consistent with the given graph $G_{n}$ harvested by contact tracing at the $n$-th stage. Even though $\mathbb{G}_{N}$ is unknown, it is reasonable for a contact tracer to assume a possible infection order in $G_{n}$ by a permitted permutation given as $\sigma=\{v_{1}=v,v_{2},...,v_{n}\}$. Thus, we can reformulate problem (1) as:
@@
where $\Omega(G_{n}\mid v)$ is the collection of all permitted permutations for $G_{n}$ rooted at $v$, and $\mathbb{P}(\sigma|v)$ is the probability of the permitted permutation $\sigma$ with $v$ as source.
Under the SI spreading model used in [15, 10, 23], the probability of a susceptible node being infected in the next time period is uniformly distributed among all outreaching edges from the infected nodes at the boundary of $G_{n}$, and the probability of a permitted permutation is as follows:
@@ II-C Regular Trees and General Graphs
From (3), we can observe that when $\mathbb{G}$ is a regular tree, both $\Phi_{i}$ and $d(v_{i})$ are constants which implies the value of (3) is also a constant. Thus, the computation of the likelihood only depends on the value of $|\Omega(G_{n},v)|$ in (2), which simplifies the computation. On the other hand, it is still unknown whether there exists an efficient method for calculating the likelihood when $\mathbb{G}$ is a general graph since $\mathbb{P}(\sigma\mid v)$ is different for each possible $\sigma\in\Omega(G_{n},v)$ and the size of $\Omega(G_{n},v)$ grows exponentially with $n$. To overcome this computational complexity issue, we aim to leverage GNN to find the MLE. Various theoretical results on regular trees or unicyclic regular graphs [10, 23] provide us a way to generate the high-quality dataset; the label of each piece of data is the exact MLE of the epidemic graph for training GNN. The analysis in Theorem 1, 2 shows the best and the worst possible outcomes during the contact tracing on regular trees, which provides us bounds on the performance of the contact tracing on general graphs.
Coming back to the contact tracer starts from the index case $G_{1}$ and collects more data in a forward manner (i.e., enlarging $G_{n}$ in (2)), the contact tracer also predicts the superspreader for that instant by solving (1). Intuitively, this means that as the contact tracing subgraph $G_{n}$ grows, the contact tracer desires this prediction to be closer (in terms of the number of hops in $\mathbb{G}_{N}$) to the most likely superspreader in $\mathbb{G}_{N}$. Considering this, how should $G_{n}$ grow in size in forward contact tracing? We propose to enlarge $G_{n}$ using BFS or DFS graph traversal [34, 35], which are detailed in the next section.